Search this course Ch 08 Assignment-Basic Stock Valuation Rabbit Builders's free cash flows (FCFs) are expected to grow at a constant long-term growth rate (g) of 1 1% per year into the future. Next year, the company expects to generate a free cash flow of $7,500,000. The market value of Red Rabit's outstanding debt and preferred stock is 56,250,000 and $31,250,000, respectively. Red Rabbit has 7,5 00,000 shares of common stock outstanding, and its weighted average cost of capital (WACC) is 17%. Given the preceding information, complete the adjacent table (rounding each value to the nearest whole doliar), and assuming that the firm has not had any nonoperating assets in its balance sheet Term Value Value of Operations es Value of Firm's Common Equity Value of Common Stock (per share) Oops, a more carefui review of Red Rabbit's balance sheet actually reports a $2,620,000 pgrtfolio of marketable securities. How does this new information affect the intrinsic value of Red Rabbit's common equity (expressed on a per-share basis) assuming no other changes to the Red Rabbir financial situation? Review the statements below and select those that accurately descrbe Red Rabbe's financial situation. Cheick all that apply The revised intrinsic value of Red Rabbit's common stock is $5.35 per share The intrinsic value of Red Rabbit's common stock decreases with the inclusion of the company's marketable securities portfolio into the analysis e intrinsic vaht of Red Rabbit's common stock İncreases wth the incusion or the company's marketable securities portfolio into t analysis
Search this course CK valuation Red Rabbit Bilders's free cash flows (FCFs) are expected to grow at a constant long-term growth rate (gu of її% per year into the future he t year the company expects to generate a free cash flow of $7,500,000. The market value of Red Rabbit's outstanding debt and preferred stock is $56,250,000 and $31,250,000, respectively. Red Rabbit has 7,500,000 shares of common stock outstanding, and its weighted average cost of capital (WACC) is 17%. information, complete the adjacent table (rounding each value to the nearest whole dollar), and assuming that the had any nonoperating assets in its balance sheet. Term Value Value of Operations Value of Firm's Common Equity Value of Comman Stock (per share) oops, a more careful review of Red Rabbit's balance sheet actually reports a $2,620,000 portfolio of marketable securities. How does this new information affect the intrinsic value of Red Rabbit's common equity (expressed on a per-share basis) assuming no other changes to the Red Rabbit financial situation? Review the statements below and select those that accurately describe Red Rabbit's financial situation. Check all that apply The revised intrinsic value of Red Rabbit's common stock is $5.35 per share. □ The intrinsic value of Red Rabbit's common stock decreases with the inclusion of the company's marketable securities portfolio into the analysis. □ The intrinsic value of Red Rabbit's common stock increases witn the inclusion of the company's marketable securtes portfol0 into the analysis The intrinsic value of the company's common stock isn't affected by the new information. CBook Pro
Answer #1

Expected FCF = FCF1 = $ 7500000, Perpetual growth rate = 11% and WACC = 17 %

Value of Operations = FCF1 / (WACC - g) = 7500000 / (0.17-0.11) = $ 125000000 or $ 125 million

Value of Debt = $ 56250000 or $ 56.25 million and Value of Preferred Stocks = $ 31250000 or $ 31.25 million

Value of Firm's Common Equity = 125 - 56.25 - 31.25 = $ 37.5 million

Number of Shares Outstanding = 7500000 or 7.5 million

Value of Common Stock (per Share) = 37.5 / 7.5 = $ 5

Presence of Marketabel Securities:

Value of Marketable Securities = $ 2.62 million

Value of Common Equity = Value of Operations - Value of Debt - Value of Preferred Stock + Value of Marketable Securities = 125 - 56.25 - 31.25 + 2.62 = $ 40.12 million

Value of Common Equity per Share = 40.12 / 7.5 = $ 5.35

As is observable, the revised intrinsic price per share increases to $ 5.35 and the overall intrinsic value of common equity also increases.Hence, the correct options are (a) and (c).